# Tracking Patterns in Toxicity and Antisocial Behavior Over User Lifetimes on Large Social Media Platforms

### v1

@@ Discussion
In summary, we’ve seen that users, on average, show different amounts of antisocial behavior depending on how long they’ve been posting. However, the direction of change depends on when they’re posting: before 2013-2015, users get less toxic over their lifetimes, while after that time, users got more toxic over their lifetimes. Also, users who would go on to post more were different from their very first post. On Reddit, high-activity users are more toxic than lower-activity users, whereas on Wikipedia, higher-activity users were less toxic. When tracking lifetimes of information instead of users, URLs that would go on to draw more discussion (higher-activity information) generally had more toxic discussion from the beginning, similar to the pattern for Reddit users.
Given that our work describes a change happening in 2013-2015, it’s useful to note other works that have observed changes around the mid-2010s. Waller and Anderson[39] find a shift in polarization of Reddit communities in 2016, and Haidt’s new book[47] discusses a “tidal wave” of adolescent mental illness around 2013, which he suggests is due to the rise of smartphones.[48] Connections between these and other potential factors may be an interesting avenue for future research.
Using the methodology in this paper, we have surfaced several related patterns in social media data: that users stratify by activity with respect to toxicity level, that this toxicity level changes over user lifetimes, and that these trends have not been stable over the history of the platform. This methodology may also be useful to form the basis for other investigations of how social media can affect user behavior.
@@ Methods
We used the Perspective API[43] “SEVERE_TOXICITY” score as our toxicity metric. The API defines “SEVERE_TOXICITY” as “A very hateful, aggressive, disrespectful comment or otherwise very likely to make a user leave a discussion or give up on sharing their perspective”,[20] and uses machine learning to assign text a score between 0 and 1 based on how well it fits that definition. Scores were generated for the text of each comment in the dataset, or in the case of Wikipedia, for the text of each talk page edit. The Perspective model is regularly updated, so scores for the same piece of text will change over time. Scores for the Reddit data were generated between August 2023 and February 2024. Scores for the Wikipedia data were included with the Wikiconv dataset,[49] and generated sometime before 2018.
@@
Reddit’s content ranking depends on user voting, where any user can “upvote” or “downvote” a post or comment. Reddit asks users to vote based on whether a comment “contributes to conversation”, and is on-topic to the community it’s posted in.[26] To prevent manipulation, Reddit only publishes the total score (upvotes minus downvotes), and some “fuzzing” is applied. By “downvote score”, we mean the percentage of comments that are downvoted in a given set of comments. The downvote score is a binary value for each Reddit comment: a comment is considered “downvoted” if it got more downvotes than upvotes, meaning its vote score is 0 or lower (since all comments start with a vote score of 1.) Vote score thresholds of -1, -2, -3, -5, and -10 were qualitatively similar; see Supplementary Section S.E. (Positive thresholds, however, were different.)
@@
Reddit data was obtained from the Pushshift archive of all comment and posts on Reddit.[50] The Reddit user sample consists of all comments made by a sample of 1 million user accounts between January 2005 and December 2022. This represents roughly 1% of the 98.6 million users who commented on Reddit during that period. Users were randomly sampled from all users who had made at least one, but less than 50,000, comments.
The 50,000-comment threshold was chosen to avoid bots and other unusual accounts after a rough analysis of high-volume users: 27% of accounts with > 100,000 comments have “bot” in their name, while only 0.25% of accounts with < 10,000 comments do. In our sample, only 1,200 users (0.12%) have more than 10,000 comments. Of those, seven have “bot” in the name, and on inspection six of those seven are actually bots.
@@
After filtering, our dataset had 123 million comments.
@@
Wikipedia analysis used the Wikiconv dataset[49] of all Wikipedia talk page edits. Bot accounts were roughly filtered out by removing any account with a name ending with “bot” (case-insensitive), as Wikipedia policy recommends for bot accounts.[51] Accounts containing “MediaWiki” in the name were also removed. After filtering, the dataset had 122 million edits from 2.8 million authors. When analyzing by pages, we ignore edits to user talk pages, which generally contain interpersonal conversations rather than discussion of a specific topic. User talk pages comprise $\sim 60\%$ of edits in the dataset; versions of Figures 2 and 4 for user talk pages can be found in Supplementary Section S.A.2.
@@
Reddit posts can either be text, images, videos, or URLs. To compare conversations about the same topic over time, we chose a subset of URL posts: YouTube and Imgur links. This choice was made since these URLs are easily canonicalized, meaning that links that point to the same page can be identified. Canonicalization means we can be more confident that we’ve collected all discussions of the same URL into one group. The post dataset consists of all posts of YouTube and Imgur URLs that were posted at least 5 times; only posts with at least one comment were counted. The dataset has 104 million comments on 11 million posts. For each post, an average score for each metric was calculated using all comments on that post.
@@
Author contributions statement
K.B. and J.K. conceived the experiment(s), K.B. analyzed the data. All authors reviewed the manuscript.
@@
Data availability All data used is publicly available.
@@
Structure is the same as Fig. 3; individual plots have the same structure as Fig. 2, with 95% confidence intervals and trend lines fitted to bucket means.
@@
The structure is the same as Fig. 2, with 95% confidence intervals and trend lines fitted to bucket means.
@@
The structure is the same as Fig. 4. The lower plots show the cumulative sum of the upper plots. The minimum value for each lifetime activity bucket is marked by a vertical line. If there were an internal minimum, that would be the point where the slopes could be said to cross zero. For visibility, the slopes are scaled so that the maximum absolute value is 1 for each lifetime activity bucket.
@@
The structure is the same as Fig. 7. The lower plots show the cumulative sum of the upper plots. The minimum value for each lifetime activity bucket is marked by a vertical line. If there were an internal minimum, that would be the point where the slopes could be said to cross zero. For visibility, the slopes are scaled so that the maximum absolute value is 1 for each lifetime activity bucket.
Appendix S.B Bootstrapping: robustness check for date of trend reversal

### v2

@@ Discussion
In summary, we’ve seen that users, on average, show different amounts of antisocial behavior depending on how long they’ve been posting. However, the direction of change over their lifetime depends on when they began posting: before 2013-2015, users get less toxic over their lifetimes, while after that time, users got more toxic over their lifetimes. Also, users who would go on to post more were different from their very first post. On Reddit, high-activity users are more toxic than lower-activity users, whereas on Wikipedia, higher-activity users were less toxic. When tracking lifetimes of information instead of users, URLs that would go on to draw more discussion (higher-activity information) generally had more toxic discussion from the beginning, similar to the pattern for Reddit users.
It is interesting to ask how these findings about the contrast in toxicity by activity level — first for Reddit, then for Wikipedia — accord with intuition about social media. Earlier, it may have seemed “obvious” that the most active users were the most toxic, given social media’s reputation for toxic behavior. When Figure 6(a) shows the opposite, it might seem equally “obvious” that people with toxic behavior by definition don’t follow norms and can’t last long. But the point is that both of these “obvious” outcomes are observed, in different settings. To quote Duncan Watts in Everything is Obvious, “When every answer and its opposite appears equally obvious […] something is wrong with the entire argument of ‘obviousness’.”[50, 51] The question is thus not which tendency is true of social media in general - whether high- or low-activity users are more toxic - but which, if any, is characteristic of a particular community. Viewed this way, as a comparison across different sites, it accords with some underlying intuition that high-activity Wikipedia users are less toxic than low-activity ones.[52] Indeed, Wikipedia is known for its strongly enforced social norms, and is often favorably compared to other social media giants; Wikipedia has been described as “one of the few remaining places that retains the faintly utopian glow of the early World Wide Web”.[48]
Given that our work describes a change happening in 2013-2015, it’s useful to note other works that have observed changes around the mid-2010s. Waller and Anderson[39] find a shift in polarization of Reddit communities in 2016, and Jonathan Haidt’s new book[53] discusses a “tidal wave” of adolescent mental illness around 2013[54], which he suggests is due to the rise of smartphones.[55] While the methods used here aren’t well suited to establishing causal mechanisms, connections between these and other potential factors may be an interesting avenue for future research.
Using the methodology in this paper, we have surfaced several related patterns in social media data: that users stratify by activity with respect to toxicity level, that this toxicity level changes over user lifetimes, and that these trends have not been stable over the history of the platform. This methodology may also be useful to form the basis for other investigations of how social media can affect user behavior. Finally, several limitations should be acknowledged. Most importantly, the analysis can’t determine the cause of the patterns we’ve found. While our findings are consistent across metrics and platforms, the multi-decade time period we analyze encompasses many changes to social norms, user demographics, technology, political polarization, platform moderation, external manipulation, and more - any of which could affect our results. The datasets also have limitations. On Reddit, 10% of comments are deleted and unavailable for analysis, which could confound results, especially if toxic comments were more likely to be deleted. Neither metric used can perfectly capture antisocial behavior - as previously discussed, the Perspective API toxicity score has known flaws and biases, and the downvote metric captures comments that are downvoted for reasons other than antisocial behavior. Our analysis is also very high-level, aggregating across large groups of users. Future work could analyze user trajectories at a more granular level, distilling patterns in the behavior of individual users and perhaps providing more insight into causality.
@@ Methods
We used the Perspective API[45] “SEVERE_TOXICITY” score as our toxicity metric. The API defines “SEVERE_TOXICITY” as “A very hateful, aggressive, disrespectful comment or otherwise very likely to make a user leave a discussion or give up on sharing their perspective”,[20] and uses machine learning to assign text a score between 0 and 1 based on how well the text fits that definition. Scores were generated for the text of each comment in the dataset, or in the case of Wikipedia, for the text of each talk page edit. The Perspective model is regularly updated, so scores for the same piece of text will change over time. Scores for the Reddit data were generated between August 2023 and February 2024. Scores for the Wikipedia data were included with the Wikiconv dataset,[56] and generated sometime before 2018.
Toxicity scores for non-English text were ignored for most datasets, due to the potential for between-language bias.[46] For the Reddit user dataset, English text was determined using Perspective API’s language detection feature. For Wikipedia, this was done by only generating scores for the English section of Wikipedia. Comments were not filtered by language for the Reddit post URL dataset, due to the additional cost of language detection for such a large number of comments.
@@
Reddit’s content ranking depends on user voting, where any user can “upvote” or “downvote” a post or comment. Reddit asks users to vote based on whether a comment “contributes to conversation”, and is on-topic to the community it’s posted in.[26] To prevent manipulation, Reddit only publishes the total score (upvotes minus downvotes), and some “fuzzing” is applied. By “downvote score”, we mean the percentage of comments that are downvoted in a given set of comments. The downvote score is a binary value for each Reddit comment: a comment is considered “downvoted” if it got more downvotes than upvotes, meaning its vote score is 0 or lower (since all comments start with a vote score of 1). Vote score thresholds of -1 and -2 were qualitatively similar; see Supplementary Section S.F. (Positive thresholds, however, were different.)
@@
Reddit data was obtained from the Pushshift archive of all comments and posts on Reddit.[57] The Reddit user sample consists of all comments made by a sample of 1 million user accounts between January 2005 and December 2022. This represents roughly 1% of the 98.6 million users who commented on Reddit during that period. Users were randomly sampled from all users who had made at least one, but fewer than 50,000 comments.
The 50,000-comment threshold was chosen to avoid bots and other unusual accounts after a rough analysis of high-volume users: 27% of accounts with > 100,000 comments have “bot” in their name, while only 0.25% of accounts with < 10,000 comments do. In our sample, only 1,200 users (0.12%) have more than 10,000 comments. Of those, seven have “bot” in the name, and on inspection, six of those seven are actually bots.
@@
After filtering, our dataset had 123 million comments (Table 1).
92.7% of comments were entirely in English, and 95.5% contained at least some English (Supplementary Section S.I). Toxicity scores for comments that were not entirely in English were ignored, as scoring accuracy could be compromised.[46] Results remained the same after this change.
Roughly 10% of comments on Reddit are deleted (Supplementary Section S.J), When comments are deleted, author information is also removed, so it’s impossible to tell which comments belong to authors in our sample. Readers should be aware that some data are missing.
@@
Wikipedia analysis used the Wikiconv dataset[56] of all talk page edits on English Wikipedia. Bot accounts were roughly filtered out by removing any account with a name ending with “bot” (case-insensitive), as Wikipedia policy recommends for bot accounts.[58] Accounts containing “MediaWiki” in the name were also removed. After filtering, the dataset had 122 million edits from 2.8 million authors (Table 1). When analyzing by pages, we ignore edits to user talk pages, which generally contain interpersonal conversations rather than discussions of a specific topic. User talk pages comprise $\sim 60\%$ of edits in the dataset; versions of Figures 2 and 4 for user talk pages can be found in Supplementary Section S.A.2.
@@
Reddit posts can either be text, images, videos, or URLs. To compare conversations about the same topic over time, we chose a subset of URL posts: YouTube and Imgur links. This choice was made since these URLs are easily canonicalized, meaning that links that point to the same page can be identified. Canonicalization means we can be more confident that we’ve collected all discussions of the same URL into one group. The post dataset consists of all posts of YouTube and Imgur URLs that were posted at least 5 times; only posts with at least one comment were counted. The dataset has 104 million comments on 11 million posts (Table 1). For each post, an average score for each metric was calculated using all comments on that post.
@@
Author contributions statement K.B. and J.K. conceived the experiment(s), K.B. analyzed the data. All authors reviewed the manuscript.
@@
Data availability The Wikipedia dataset analysed during the current study is available in the Wikiconv repository, https://github.com/conversationai/wikidetox/tree/main/wikiconv. The Reddit dataset analysed during the current study was publicly available from Pushshift at the time of analysis, but Pushshift data is currently available only under the request procedure outlined at https://web.archive.org/web/20250327162604/https://www.reddit.com/r/pushshift/comments/14ei799/pushshift_live_again_and_how_moderators_can/.
Supplementary Material
@@
Structure is the same as Fig. 3; individual plots have the same structure as Fig. 2, showing 95% confidence intervals and trend lines fitted to bucket means.
@@
The structure is the same as Fig. 2, showing 95% confidence intervals and trend lines fitted to bucket means. Asterisks represent confidence levels for each slope.
@@
The structure is the same as Fig. 4/Fig.S15. The lower plots show the cumulative sum of the upper plots. The minimum value for each lifetime activity bucket is marked by a vertical line. If there were an internal minimum, that would be the point where the slopes could be said to cross zero. For visibility, the slopes are scaled so that the maximum absolute value is 1 for each lifetime activity bucket.
@@
The structure is the same as Fig. 7/Fig.S16. The lower plots show the cumulative sum of the upper plots. The minimum value for each lifetime activity bucket is marked by a vertical line. If there were an internal minimum, that would be the point where the slopes could be said to cross zero. For visibility, the slopes are scaled so that the maximum absolute value is 1 for each lifetime activity bucket.
Appendix S.B Statistical validation
S.B.1 Bootstrapping: robustness check for date of trend reversal